Circular No. 6122

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PERIODIC COMET SCHWASSMANN-WACHMANN 3 (1994w) K. Birkle, Max Planck Institute for Astronomy, Heidelberg; H. Böhn-

hardt, Astronomical Institute, University of Munich; and G. Schwehm, ES-TEC, Noordwijk, report their recovery of this comet with the prime-focus CCD camera on the 3.5-m reflector at Calar Alto, as follows:

1994 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 28.03183  10^{h}29^{m}24.^{s}2  +23^{o}50′51′′  22
29.07865  10 29 22.7  +23 57 07

There was no detectable coma. The indicated correction to the prediction on MPC 20123 is ΔT = +0.15 day.

SUPERNOVA 1994aj IN ANONYMOUS GALAXY S. Benetti, European Southern Observatory (ESO), reports: “A fully-

reduced CCD spectrum (range 450–700 nm, resolution 1.2 nm), obtained on Dec. 31.22 UT with the ESO 2.2-m telescope (+ EFOSC2), of SN 1994aj (cf. IAUC 6121) shows Hα and Hβ lines with a complex shape superimposed on a relatively blue continuum. The Balmer lines consist of a broad compoment (FWHM ∼ 7000 km/s) in emission, superimposed by a second component having a narrow P-Cyg profile (expansion velocity 1000 km/s and emission peaks measured at 677.5 and 501.6 nm). Other lines present in the spectrum, such as Fe II and Na I D, seem to have normal P-Cyg profiles. The expansion velocity derived from the minimum of the Na I D lines, corrected by the recession velocity of 9600 km/s of the parent galaxy (measured from a faint Hα emission present in the galaxy spectrum) is ∼ 3500 km/s. Spectroscopic and photometric observations are greatly needed to follow the evolution of this unique type-II supernova.”

POSSIBLE SUPERNOVA IN NGC 2782 S. Nakano, Sumoto, Japan, reports the discovery of a possible super-

nova (mag 16.3) by Reiki Kushida at Yatsugatake South Base Observatory. Found on an unfiltered CCD image taken on Dec. 24.856 UT, the candidate’s position is α = 9^{h}10^{m}50^{s}.13, δ = +40^{o}18′47′′.1 (equinox 1950.0; measured by Y. Kushida), which is 50′′ west and 30′′ south of the center of NGC 2782. Additional CCD magnitudes: Dec. 17, [16.4; 27.762, 16.1. A star of mag 16.7 is located 60′′ west and 35′′ north of the galaxy’s center.

1994 December 31 Daniel W. E. Green